# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v2

@@
Abstract
The wide range of sub- and super-nuclear densities achieved in neutron stars makes them ideal probes of dense nuclear behavior in the form of the nuclear equation of state (EoS). Studying neutron stars both in isolation and in highly dynamic events, many recent observations, most famously the gravitational wave and electromagnetic signals associated with the BNS merger GW170817/AT2017gfo, have provided suggestive insight into these highest nuclear densities. Measurements of galactic neutron star masses and radii from NICER and other radio and X-ray measurements provide critical complementary perspectives, bounding other features of the EoS. Though nominally congruent, in this paper we highlight many underappreciated “hidden” priors embedded in joint analysis of these many messengers, and their systematic impact on joint EoS inference. In this work, we perform a careful step-by-step Bayesian inference using a simple low-dimensional parametric EoS model, incrementally adding information from galactic pulsars, gravitational wave sources, and more speculative constraints involving kilonovae and the neutron star maximum mass. At each stage, we carefully discuss the marginal likelihood, explaining how hidden priors impact conclusions. Conversely, we also quantify how much information these measurements have, arguing that many if not most have minimal impact relative to the explicit and hidden prior assumptions. Specifically, we find two features dominate our inference: on the one hand, the choice of EoS parameterization, and on the other the astrophysical priors associated with interpreting events, particularly the multimessenger source GW170817. Spectral parameterizations have been shown to cause large differences in the predicted EoS based on if causality is applied in the framework or not, and applying causality predicts a strong phase-transition-like feature at $\sim 10^{14.5}$ g cm-3. In an appendix, we outline a simple semianalytic projection suitable for assessing the measurability of the EoS with ensembles of present and future detections.
I Introduction
@@
Neutron star observations, made over the last decades have provided a critical perspective into matter at high densities, complementing rapid theoretical and experimental progress [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 5, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. High-precision pulsar observations have provided evidence for a wide range of neutron star masses, constraining the neutron star maximum mass from below [33]. X-ray measurements of neutron star radii have provided additional constraints [32, 34, 35, 36]. Recently, the NICER mission has enabled unique neutron star periodic profile monitoring, enabling simultaneous mass and radius constraints on two neutron stars [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Observations of merging neutron stars have also provided several signatures which potentially limit the nuclear equation of state (EoS). Gravitational wave measurements of GW170817 limit the effective inspiral tidal deformability [3, 12, 52, 14, 53, 54, 24]. Simultaneous electromagnetic observations also provide suggestive hints into neutron star compactness, deduced from the presence and energetics of gamma-ray [55] and kilonova [28, 56, 57] emission [3, 12, 13, 14, 58, 59, 22, 60, 61]. While the origin of quasi-periodic oscillations remain unclear, if these can be attributed to neutron stars, then additional constraints can be formulated [62].
The different approaches largely suggest a comparable neutron star radius (see, e.g., [32] and references therein), as deduced by many groups employing similar Bayesian methods from various independent observational sources and techniques (see Ref. [63] for a review of Bayesian statistics for astrophysical observations). However, the individual underlying inferences involve many often subtle steps which disguise their implicit and explicit priors, particularly when combining very heterogeneous sources of external information. Unsurprisingly and closely related, looking more closely at individual measurements and their systematics suggests tension between these different observational approaches to the EoS. For example, several studies have pointed out that the amount and nature of material ejected from a merger like GW170817 is inconsistent with the matter needed to reproduce the details of the kilonova emission [64, 65, 66, 67, 68]. Conversely, several investigations have wholeheartedly embraced combining all available measurements to produce overall estimates, despite the incremental impact most of these measurements have on overall inferences, owing to large statistical measurement errors (let alone poorly understood systematics); see, e.g., the discussion in [28]. Finally, as highlighted by previous work [69, 32, 70, 71], the relative impact of all measurements derives strongly from their explicit prior assumptions, both on the small and large scale. For example, compared to theoretically well-motivated priors [72, 73], adopting an extremely conservative phenomenological approach with much wider priors enables individual measurements to strongly rule out individual realizations (e.g., [74]). Conversely, a too-tightly-parameterized approach could conceivably miss subtle degeneracies associated with putative phase transitions; see, e.g., [75] and references therein.
@@
In this work, we revisit how to combine several different constraints on the nuclear EoS, including explicit and implicit theoretical or astrophysical priors. We demonstrate these inferences rely on rarely discussed and of necessity strong assumptions about the neutron star population. We argue these often inconsistently-applied prior factors have a substantial effect and may cause analyses to inadvertently rule out a priori consistent EoS models. Conversely, we demonstrate that most extant measurements are largely consistent, with modest tension with each other and our priors even for our most extreme scenarios.
Unlike many recent studies which rely on increasingly complex models for the nuclear EoS, to preserve our emphasis on the observational ingredients and priors we employ an extremely simple low-dimensional model for the nuclear EoS:
@@
two conventional spectral representations [76, 77, 78] based on approximating a functional versus (log) pressure with a Taylor series. These low-dimensional smooth approximations are known to have limited flexibility compared with more expressive infinite-dimensional nonparametric models like Gaussian processes [21, 5, 79]. However, these models are more than sufficient to express contemporary observational uncertainties, helping us indicate avenues where systematics may play an important role. Other studies have thoroughly explored more flexible approaches including multiple phase transitions; see, e..g, [74] and references therein.
@@
This paper is organized as follows. In Section II we review our phenomenological spectral approach to the nuclear equation of state; pertinent observations of neutron stars; our specific implementation of hierarchical Bayesian inference for the nuclear equation of state; and our specific procedure and priors for each class of measurement. For galactic neutron stars we employ radio observations of heavy neutron stars PSR-J0348 [80] and PSR-J1641 [81, 82], NICER X-ray pulse profiling for PSR-J0030, PSR-J0740, and X-ray spectra and distance analyzed for HESS J1731 [83] which retains the possibility of consisting of strange quark matter, however for this study we assume it to be nucleonic. For gravitational wave sources, we employ only GW170817. In Section III we provide inferences for the nuclear EoS using galactic pulsars and nuclear symmetry energy; update that inference using information from GW170817, and then reassess those results by incorporating more indirect interpretations of observations, including limits on the NS maximum mass. Finally, we present our concluding remarks in Sec. IV.
II Methods
II.1 Nuclear equation of state phenomenology

### v3

# Astrophysical assumptions and equation of state framework have larger impact on equation of state inference than individual neutron star observations

@@
11institutetext: Center for Computational Relativity and Gravitation, Rochester Institute of Technology, Rochester, New York 14623, USA22institutetext: Department of Physics, North Carolina State University, Raleigh, North Carolina 27695, USA33institutetext: Department of Physics, Kenyon College, Gambier, Ohio 43022, USA
Astrophysical assumptions and equation of state framework have larger impact on equation of state inference than individual neutron star observations
Atul Kedia    Richard O’Shaughnessy    Leslie Wade    Anjali Yelikar
The wide range of nuclear densities achieved in neutron stars makes them probes of dense nuclear behavior in the form of the nuclear equation of state (EoS). Studying neutron stars both in isolation, with X-ray measurements and pulse profiling, and in dynamic events, such as neutron star mergers, have provided insight into these high nuclear densities. Though nominally congruent, here we highlight impact of implicit assumptions embedded in joint analysis of these messengers and their systematic impact on EoS inference. We show that astrophysical assumptions and EoS framework can have a larger effect on inferred EoS than individual contemporary neutron star observations. Performing a proof-of-concept demonstration using the chronologically first few observational constraints, after the application of 5 to 6 observational constraints, additional observations provided diminishing returns and modified the inferred EoS by shifting the radius of a 1.4 $\rm M_{\odot}$} NS by $\sim$ 0.1 km. By contrast, astrophysical priors, specifically the spin and mass ratio motivated by astrophysical source population uncertainties, and EoS framework tend to impact EoS inference much more substantially by shifting the 1.4 $\rm M_{\odot}$} NS radius by $\sim$ 0.5 km and by modifying shape of inferred mass-radius relationship. The inferred EoS depends strongly on the adopted choice of spectral parameterizations: when we employ a framework which explicitly enforces causality, we find a strong phase-transition-like feature at $\sim 10^{14.5}$ g cm-3.
1 Introduction
@@
Neutron star observations, made over the last decades have provided a critical perspective into matter at high densities, complementing rapid theoretical and experimental progress (LIGO-GW170817-bns; LIGO-GW170817-SourceProperties; LIGO-GW170817-EOS; 2020GReGr..52..109C; Landry:2020vaw; 2023PhRvD.107d3034M; Reed:2021nqk; Essick:2021kjb; Agathos:2015uaa; Lackey:2014fwa; Alvarez-Castillo:2016oln; Margalit:2017dij; Annala:2017llu; Most:2018hfd; Radice:2017lry; Rezzolla:2017aly; Riley:2018ekf; Tews:2018iwm; Greif:2018njt; Kiuchi:2019lls; Landry:2020vaw; Landry:2018prd; Shibata:2019ctb; 2020NatAs...4..625C; 2020PhRvR...2c3514P; 2020ApJ...893...61Z; Essick:2020flb; 2020ApJ...893L..21R; Dietrich:2020efo; 2020EPJST.229.3663C; 2020PhRvC.102d5807L; 2021PhRvC.103c5802X; 2021PhRvL.126f1101A). High-precision pulsar observations have provided evidence for a wide range of neutron star masses, constraining the neutron star maximum mass from below (2016ARA&A..54..401O). X-ray measurements of neutron star radii have provided additional constraints (2021PhRvL.126f1101A; 2016ApJ...831..184B; Steiner:2017vmg; 2023JCAP...02..016F). The NICER mission has enabled unique neutron star periodic profile monitoring, enabling simultaneous mass and radius constraints on a handful of neutron stars including J0740, J0437, and J0614 (Miller:2019cac; Riley:2019yda; Raaijmakers:2019qny; Bilous:2019knh; 2019ApJ...887L..27G; Bogdanov:2019ixe; Bogdanov:2019qjb; 2021ApJ...914L..15B; Miller:2021qha; Riley:2021pdl; 2021ApJ...918L..29R; 2020ApJ...894L...8C; Raaijmakers:2019dks; Legred:2021; Pang:2021jta; 2024ApJ...976...58S; 2025arXiv250614883M; 2024ApJ...971L..20C) Observations of merging neutron stars have also provided several signatures which potentially limit the nuclear equation of state (EoS). Gravitational wave measurements of GW170817 limit the effective inspiral tidal deformability (LIGO-GW170817-EOS; Margalit:2017dij; De:2018uhw; Most:2018hfd; 2018ApJ...852L..29R; 2017ApJ...850L..34B; 2020PhRvR...2c3514P). Simultaneous electromagnetic observations also provide suggestive hints into neutron star compactness, deduced from the presence and energetics of gamma-ray (2018ApJ...852L..25R) and kilonova (Dietrich:2020efo; Coughlin:2018miv; Coughlin:2018fis) emission (LIGO-GW170817-EOS; Margalit:2017dij; Annala:2017llu; Most:2018hfd; Raithel:2018ncd; 2018PhRvD..97b1501R; Shibata:2019ctb; Most:2018eaw; 2021ApJ...908L..28N). While the origin of quasi-periodic oscillations remain unclear, if these can be attributed to neutron stars, then additional constraints can be formulated (2023A&A...676A..65S).
Initial interpretations of the first GW and two NICER measurements suggested a consensus towards a comparable NS radius (see, e.g., 2021PhRvL.126f1101A and references therein), as deduced by many groups employing similar Bayesian methods from various independent observational sources and techniques (see 2019PASA...36...10T for a review of Bayesian statistics for astrophysical observations). Even after modest revisions with updated modeling systematics, the original consensus remained intact; see, e.g., 2024ApJ...971L..19R; 2025PhLB..86539501L; 2025PhRvX..15b1014K and references therein. However, at the time of the the most recent observation of J0614 (2025arXiv250614883M) , the full range of the handful of NICER observations are individually consistent with a broad range of NS radii, particularly once measurement systematis are included (Riley:2019yda; 2019ApJ...887L..27G; 2024ApJ...961...62V). Similarly, intial multimessenger inferences of GW and the kilonova associated with AT2017gfo suggested the light curves provided modest additional information further supporting the initial consensus (Dietrich:2020efo). Since then, however, several studies have pointed out that the amount and nature of material ejected from a merger like GW170817 is currently inconsistent with this consensus and our present understanding of kilonova emission (2021ApJ...906...98N; 2020ARNPS..7013120R; 2021ApJ...910..116K; PhysRevResearch.5.013168; Ristic22; 2025arXiv250312320R). In sum, preliminary conclusions about the nuclear EOS based early multimessenger evidence are narrower than what now seems appropriate given the latest observations and understanding of instrumental and modeling systematics.
@@
In this work, motivated by the apparent tension between early estimates and present observations, we revisit the observations used to establish the initial consensus, to seek out underappreciated or unknown systematic factors in the data analysis strategy which can significantly impact the inferred EOS. Using an independent Bayesian framework to combine multiple independent observations, we demonstrate these inferences rely on rarely discussed and of necessity strong assumptions about the neutron star populations pertinent to each class of observation. We argue these often inconsistently-applied prior factors have a substantial effect and may cause analyses to inadvertently rule out a priori consistent EoS models. Conversely, we demonstrate that these measurements are largely consistent, with modest tension with each other and our priors even for our most extreme scenarios.
@@
Unlike many recent studies which rely on increasingly complex models for the nuclear EoS, to preserve our emphasis on the observational ingredients and priors we employ an extremely simple low-dimensional model for the nuclear EoS: two conventional spectral representations (Lindblom:2010bb; Lindblom:2018rfr; PhysRevD.105.063031) based on approximating a functional versus (log) pressure with a Taylor series. These low-dimensional smooth approximations are known to have limited flexibility compared with more expressive infinite-dimensional nonparametric models like Gaussian processes (Landry:2018prd; Landry:2020vaw; Landry:2021hvl). However, these models are more than sufficient to express contemporary observational uncertainties, helping us indicate avenues where systematics may play an important role. Other studies have thoroughly explored more flexible approaches including multiple phase transitions; see, e..g, (PhysRevD.110.123009) and references therein.
@@
This paper is organized as follows. In Section 2 we review our specific implementation of hierarchical Bayesian inference for the nuclear equation of state. Specifically, in Sections 2.2, 2.4, 2.6 we will highlight how the likelihood has factors which depend explicitly on astrophysical assumptions about the source population. We will then discuss a range of possible source populations, particularly for the exceptional source GW170817. In this section we also review pertinent observations of neutron stars; our specific procedure and priors for each class of measurement, and our phenomenological parameterizations of the nuclear equation of state. For galactic neutron stars we employ radio observations of heavy neutron stars PSR-J0348 (Antoniadis:2013pzd) and PSR-J1641 (2018ApJS..235...37A; 2023ApJ...951L...9A), NICER X-ray pulse profiling for PSR-J0030, PSR-J0740, and X-ray spectra and distance analyzed for HESS J1731 (2022NatAs...6.1444D) which retains the possibility of consisting of strange quark matter, however for this study we assume it to be nucleonic. For gravitational wave sources, we employ only GW170817. In Section 3 we provide inferences for the nuclear EoS using galactic pulsars and nuclear symmetry energy; update that inference using information from GW170817, and then reassess those results by incorporating more indirect interpretations of observations, including limits on the NS maximum mass. In Section 4 we elaborate on novel, unfamiliar, or unexpected features of our analysis and their broader implications, focusing on how our choces of NS population models and representation of the EOS posterior differ from other work. Finally, we present our concluding remarks in Sec. 5.
2 Methods
In this section, we carefully review hierarchical Bayesian inference for the nuclear EOS, to indicate how assumptions about the astrophysical population of NS responsible for different phenomena, particularly the unique binary multimessenger source GW170817, must be incorporated into our calculations. To assess the relative significance of the binary NS population model, we compare its impact to two other well-studied factors, as enumerated in Table 1: (a) the inclusion or omission of several other astrophysical constraints, and (b) the use of an alternative EOS parameterization and prior. (In our Appendicies, we futher investigate additional even more subdominant factors, such as the impact of systematic uncertainty in the GW signal model.)
2.1 Bayesian framework
At root, Bayesian inference expresses a posterior as a likelihood times a prior (modulo a normalization). Because independent observations of any event have independent statistical uncertainties and are associated with disctinct astrophysical phenomena, we combine independent observations by multiplying their likelihoods together. Thus, if $\cal Y$} characterizes EOS parameters having a prior $p({\cal Y})$, then the likelihood used in this work will have a form
where the factors in this expression are briefly defined below and elaborated in subsequent sections. In this expression, $n$ indexes the neutron star mass radius observations and $Z_{k}$ is computed according to Eq. (4); $k$ indexes the pulsar mass measurements providing a useful constraint on the maximum neutron star mass, with ${\cal L}_{k}$ computed using one of the factors in Eq. (2); and $g$ indexes the gravitational wave observations, with the GW likelihood computed via Eq. (6). For brevity, we may use a subscript or argument $\alpha$ to refer to an expression conditioned on a specific EOS realization ${\cal Y}_{\alpha}$. In this expression, almost all factors represent marginal likelihoods and depend on assumptions about the isolated or binary NS population responsible for those events, as described in the subsections below which describe each factor in detail.
In the equation presented above, for simplicity we have not incorporated two factors superfluous to our study but essential to real analyses of the unknown neutron star populations using a large census of NS observations: survey selection biases and models for the unknown neutron star population. For example, unbiased estimates of the nuclear equation of state require self-consistent estimates of the binary neutron star population (2020arXiv200101747W). In turn, self-consistent population estimates must correctly account for parameter-dependent selection effects effects (2019MNRAS.486.1086M; 2019PhRvD.100d3012W). To simplify our discussion and focus on the systematic effects that different astrophysical assumptions about the binary NS population can introduce, we perform a proof-of-concept calculation using only a handful of well-motivated population assumptions, rather than a broad parameterized model family.
Table 1: The astrophysical and experimental constraints applied and sections and figures in which the impact of respective constraints are discussed.
Section Figure EoS Constraint Notes
GW170817 J0348 J1614 J0740 J0030 PREX J1731 J0437 $M_{max}$
3.1 5 $\Gamma$ - - - - Base NS inference
3.1 6 $\Upsilon$ - - - - EoS framework impact
3.2.1 7, 8, 9 - - - - - - - - - NS prior impact
3.2.2 10 $\Gamma$ - - - GW+PSR+PREX joint inference
3.2.2 12 $\Gamma$ - - - NS prior impact in joint inference
3.3 14 $\Gamma$ - - - Upper limit on $M_{max}$
3.4 15 $\Gamma$ - - - Impact of J1731
App. C 17 $\Gamma$ - - - Impact of waveform models
App. E 19 $\Gamma$ - - - Impact of J0437
2.2 NS with known extremal masses
Most neutron star mass measurements provide information only about the empirical pulsar star mass distribution, rather than any direct insight into the nuclear equation of state. However, a few very massive pulsars’ existence provides useful insight into the neutron star maximum mass, bounding it from below: J0740 (2021ApJ...915L..12F; Cromartie:2019kug), with pulsar mass consistent with a Gaussian with mean and standard deviation $2.14\penalty 10000\ {\rm M_{\odot}}$ and $0.1\penalty 10000\ {\rm M_{\odot}}$; J0348+0432 (2013Sci...340..448A), with pulsar mass $2.01\penalty 10000\ {\rm M_{\odot}}\pm 0.04\penalty 10000\ {\rm M_{\odot}}$; and J1614 (2010Natur.467.1081D), with pulsar mass $1.908\penalty 10000\ {\rm M_{\odot}}\pm 0.016\penalty 10000\ {\rm M_{\odot}}$. We follow (Dietrich:2020efo) to express the likelihood of an EoS consistent with these observations as the product of standard normal distributions:
where $\Phi(z)$ is the cumulative distribution of the standard normal distribution for each maximum mass measurement and are shown in Fig. 1 as horizontal bands. Using these expressions and as illustrated concretely with plots of the pertinent single-event likelihoods in Section 3.1, these NS observations strongly suggest a familiar conclusion: that the maximum NS mass is larger than $2.1M_{\odot}$. We adopt this empirical limit to define a broad NS population.
2.3 Neutron star population model 1: galactic pulsars
Deducing the equation of state from observations of neutron star properties always requires some model for the neutron star population: empirical observables limit but do not uniquely determine the neutron star mass, spin, radius, and tidal deformability. In this work, unless otherwise noted and particularly for galactic pulsars, we will assume the isolated (and binary) neutron star population is uniformly distributed in gravitational mass $m$ between fixed, EoS-independent limits $m\in[m_{min},m_{max}]$. We specifically adopt $m_{min}=0.4\penalty 10000\ {\rm M_{\odot}}$ and $m_{max}=2.1\penalty 10000\ {\rm M_{\odot}}$, with the lower and upper mass limits chosen because of the collection of NSs we will use to constrain the nuclear EoS. Similarly, unless otherwise noted, we will assume the (aligned component) of neutron star spin is distributed isotropically in a sphere and uniformly in magnitude, with a maximum dimensionless value $a=S/m^{2}<0.05$. This fiducial population closely resembles the prior assumptions used when inferring the nature of low-mass gravitational wave sources. We emphasize that our prior represents our assumptions about the astrophysical distribution of neutron stars as realized in nature, which need not reach the maximum mass allowed by the nuclear equation of state.
Aside from its EOS-independent upper limit, this fiducial NS mass distribution has one other unconventional choice: its lower limit. Looking ahead to analyses including HESS J1731, this choice allows us to incorporate the most observationally likely parameters, avoiding baking in strong prior assumptions into our analysis about the plausibility of subsolar-mass NS. For analyses omitting the HESS source, the lower limit has a negligible impact, only impacting our overall normalization.
2.4 Neutron star mass-radius constraints
NICER observations of periodic X-ray emission on millisecond pulsars provides information about their masses and radii. Though their data product is reported as a weighted posterior distribution, we find these results can be well approximated by a simple Gaussian distribution, as illustrated in Figure 1 for the pulsar PSR-J0030 (red). In Appendix A we describe the best-fitting gaussian to each NICER observations alongwith HESS-J1731, and the specific method we used to deduce its parameters. In terms of this Gaussian likelihood $L(M,R)$, we can evaluate the marginal likelihood for each equation of state $\alpha$ according to (gwastro-PENR-RIFT; LIGO-GW170817-EOSrank; 2019PASA...36...10T)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \delta(M-M_{\alpha}(s))\mbox{d}M\mbox{d}R\bigg]p(s)\mbox{d}s$
$=\int\mathcal{L}(M(s),R_{\alpha}(M(s)))\penalty 10000\ \penalty 10000\ p(s)\frac{\mbox{d}M}{\mbox{d}M_{\alpha}/\mbox{d}s}$
$={\cal F}\int\mathcal{L}(M(s),R_{\alpha}(M(s)))\penalty 10000\ \penalty 10000\ \mbox{d}M$ (4)
[…]
2.5 Neutron star population models 2: Merging GW sources
Most compact object formation models suggest that the mass and spin distribution of NS in merging binaries will differ substantially from the mass and spin distribution of isolated or accreting NS. Addtionally, possibly-strong selection biases on binary, GW, and especially multimessenger sources can further modify the appropriate population model needed for EOS inference. For this reason, we employ five distinct mass and spin distributions for binary NS sources. Unless otherwise noted and following convention for GW sources, we assume a maximum dimensionless spin of $|\chi_{i}|<0.05$ for each NS, with $\chi_{i,z}$ distributed consistent with a uniform distribution of spin magnitudes and an isotropic distribution of spin directions according to $p_{\chi}(\chi_{i,z}|\chi_{\rm max})$. Our three primary mass and spin distributions can then be enumerated as follows
  • (a) our fiducual base model, in which each NS is drawn independently from a uniform mass distribution $p_{0}(m_{1},m_{2},\chi_{1,z}\chi_{2,z})=2{\cal F}^{2}I(m_{1})I(m_{2})\Theta(m_{1}-m_{2})p_{\chi}(\chi_{1,z})p_{\chi}(\chi_{2,z})$ where $I$ is nonzero for $m\in[m_{\rm min},m_{\rm max}]$ and where the Heavyside step function $\Theta$ ensures that $m_{1}>m_{2}$];
  • (b) a distribution which requires highly unequal masses and allows for large spin $p_{as}(m_{1},m_{2},\chi_{1,z},\chi_{2,z})\propto p_{0}$ subject to $m_{2}/m_{1}\leq 0.6$ and using $\chi_{\rm max}=0.5$]; (At GW170817’s source-frame chirp mass ${\cal M}_{c}=1.186\penalty 10000\ {\rm M_{\odot}}$, this constraint requires more extreme choices for source-frame primary and secondary mass: $m_{2}<1.062\penalty 10000\ {\rm M_{\odot}}$ and $m_{1}>1.77\penalty 10000\ {\rm M_{\odot}}$.)
  • and (c) a distribution which both favors symmetric masses and requires preferentially aligned spins [i.e., in mass adoping $p_{0}$ but requiring $\chi_{i,z}>0$ and $m_{2}/m_{1}>0.9$], with the specific choice of mass ratio cutoff motivated by the range expected for GW-detected BNS sources derived from the galactic population (2020ApJ...900L..41A).
Operationally, all of these astrophysical prior choices can be implemented as alternative options for the upper and lower bounds adopted for $q,\chi_{i,z}$, when analyzing GW170817 with a nonprecessing GW model. Several sources of external information provide good reason to adopt alternative informed priors. In the discussion below, we elaborate on our motivation behind these three choices.
The third distribution (c) is strongly motivated by galactic observations and conventional field formation scenarios. These scenarios strongly favor nearly equal masses (2020ApJ...900L..41A; 2020A&A...639A.123K; Farrow:2019xnc; 2018MNRAS.481.4009V; 2020PASA...37...38V; 2020ApJ...900L..41A; 2020A&A...639A.123K). and also usually rule out large or preferentially negatively-oriented spins (2000ApJ...541..319K). (Of course, the limited statistics available from the GW census are consistent with a much broader mass ratio distribution (Landry:2021hvl; LIGO-O3-O3bpop) and provide minimal constraints on plausible NS spin orientations; additionally, non-merging galactic BNS binaries have a much broader mass ratio distribution (2023MNRAS.521.4669Y).)
The mass distribution adopted for (b) is motivated by the unexpectedly large dynamical ejecta inferred for AT2017gfo.As most predictions for this ejecta are suppressed for $q\simeq 1$ (Henkel:2022naw), others have suggested favoring asymmetric masses where the dynamical ejecta isn’t suppressed (2018ApJ...866...60P; 2020Natur.583..211F). For example, 2025arXiv250312320R argue that, if contemporary ejecta models for both the mass and velocity are applied directly, a self-consistent interpretation of AT2017gfo either requires implausible NS radii or extremely asymmetric binary mergers.
As will be demonstrated later, due to strong spin-mass ratio correlations, this constraint also requires substantial NS spin. To avoid adopting overconstraining prior assumptions, for this analysis we will substantially relax the maximum spin allowed. To better understand the impact of relaxing the maximum spin, we will also present a handful of results using the fiducial model but with a larger $\chi_{\rm max}$. This relaxed spin prior can also be well-justified astrophysically on similar grounds:others have demonstrated that a strongly-spinning secondary neutron star ($m_{2}$) will produce substantially enhanced ejecta, as its outer layers are less bound (2024MNRAS.530.2336R; 2022MNRAS.513.3646P). Strongly spinning NS mergers can occur for example if the secondary does not have time to spin down, due to fortuitous NS kicks (2010ApJ...716..615O) or Kozai-Lidov effects in hierarchical triples (2023arXiv230210350B). (Such a rapid formation scenario isn’t consistent with the large geometric offset for AT2017gfo but could occur in general; as noted in (2023arXiv230210350B; 2024arXiv240203696C) this rapid formation scenario would reconcile NS merger formation with r-process-enriched but metal-poor stars.) While the specific form adopted for the prior (b) is somewhat ad hoc, particularly given the systematics associated with contemporary ejecta calculations, an exhaustive quantitative investigation needed to develop and calibrate another approximation merits its own dedicated study.
2.6 Gravitational wave observations
Gravitational wave observations of GW170817 constrain the binary masses $m_{i}$ and tidal deformabilities $\Lambda_{i}$ for each component of the binary (LIGO-GW170817-bns; LIGO-GW170817-SourceProperties; LIGO-GW170817-EOS; PhysRevResearch.2.043039). Specifically, the presence of matter at leading order impacts the binary’s inspiral, through each components’ dimensionless tidal deformability parameter $\Lambda=(2/3)k_{2}(c^{2}R/GM)^{5}$ (Flanagan:2007ix), where $k_{2}$ is the $l=2$ Love number, $M$ is the mass, and $R$ is the radius. The leading order contribution to the phase evolution of an GW inspiral is given by the weighted combination of $\Lambda$ terms
The impact of tidal effects are included in many conventional approximate phenomenological estimates of outgoing radiation from merging compact binaries. In this work, we will focus on analyses with a contemporary state-of-the-art model, IMRPhenomPv2_NRTidalv2 (2019PhRvD..99b4029D; 2019PhRvD.100d4003D), which incorporates precession physics but omits higher-order modes. We note other contemporary models like NRHybSur3dq8Tidal (2020PhRvD.102b4031B) and TEOBResumS (2018PhRvD..98j4052N) include these higher-order modes and tidal effects. In Appendix C we present a single proof-of-concept inference using one of these alternative waveform models, to demonstrate its effect is subdominant to the issues emphasized in this work.
When performing phenomenolgical inference for binary intrinsic parameters $X=({\cal M}_{c},\eta,\mathbf{\chi}_{1},\mathbf{\chi}_{2},\Lambda_{1},\Lambda_{2})$, the RIFT parameter inference engine (gwastro-PENR-RIFT; 2023PhRvD.107b4040W) iteratively assesses the gravitational wave marginal likelihood $L(X)$ on many candidate points $X$, producing estimates of the posterior $\propto L(X)p_{gw}(X)$ for $X$ using (fiducial) population priors for these intrinsic parameters $p_{gw}(X)$ and an interpolated estimate $\hat{L}(X)$ of the true likelihood $L(X)$ obtained by interpolation. For our application, the RIFT marginal likelihood data and interpolation provide a framework to compute an evidence for any proposed EOS (gwastro-PENR-RIFT; LIGO-GW170817-EOSrank; 2024arXiv240715753V). Specifically, using the marginal likelihood $L(X)$ obtained by interpolating the RIFT marginal likelihoods and the EoS-derived relationship connecting $\Lambda(m|{\cal Y})$ to the neutron star gravitational mass $m$ and equation of state model $\cal Y$}, we compute a marginal evidence
${\cal E}=\int L(X(x,{\cal Y}))p(x)dx$ (6)
where $x=(m_{1},m_{2},\mathbf{\chi}_{1},\mathbf{\chi}_{2})$ are binary parameters omitting tides and $X(x,{\cal Y})$ supplements these parameters with the EoS-derived tidal parameters (i.e., $X=({\cal M}_{c},\eta,\mathbf{\chi}_{1},\mathbf{\chi}_{2},\Lambda(m_{1}|{\cal Y}),\Lambda(m_{2}|{\cal Y}))$). For simplicity, we also use RIFT to perform this integral (gwastro-PENR-RIFT; LIGO-GW170817-EOSrank; gwastro-RIFT_FinerNet), using the RIFT marginal likelihood estimator $\hat{L}(X)$ for $L(X)$. These EoS evidence infrastructure has been validated in previous work (LIGO-GW170817-EOSrank). We do not employ any customary approximations when evaluating this integral numerically (e.g., that the chirp mass is well known).
While other objects putatively containing matter have been identified via their GW signature (i.e., GW190425 (2020ApJ...892L...3A), GW200105, and GW200115 (LIGO-O3-NSBH)), their GW signature also does not meaningfully contain the nature of nuclear matter. For example, the two objects in NSBH binaries have putative matter effects that minimally impact $\tilde{\Lambda}$. Conversely, in GW190425, the large NS masses imply very small tidal deformabilities, but that low-amplitude GW measurement only weakly constraints $\tilde{\Lambda}$. For this reason, we omit further discussion of these three objects.
Phenomenology of $\Lambda,\penalty 10000\ R$
To provide a preliminary estimate of the impact of the NS population priors on EoS inference derived from GW170817, we follow Zhao and Lattimer 2018 (Zhao:2018nyf), and adopt simple approximate relationships between tidal deformability and fiducial NS radius $R_{1.4}$. In brief, they argue that to a good approximation $\Lambda\propto(R/M)^{6}$ (Eq. (13) in Zhao:2018nyf); that therefore $\tilde{\Lambda}$ is to a good approximation proportional to $(R_{1.4}/{\cal M}_{c})^{6}$, with a nearly mass-ratio-independent constant (Eq. (15) in Zhao:2018nyf); and therefore for a binary with well-measured chirp mass, the radius of a fiducial NS can be estimated as (evaluated for GW170817)
$R_{1.4}\simeq 13.4\penalty 10000\ {\rm km}(\tilde{\Lambda}/800)^{1/6}\;.$ (7)
2.7 Nuclear equation of state phenomenology